OpenStudy (anonymous):

i am in seventh grade...can someone show me how to solve 2 2/3 + -1 1/2= thank u

OpenStudy (anonymous):

You want to first change them both into fractions: \[2 \frac{2}{3} + (-1\frac{1}{2})\]\[\frac{8}{3} + (-\frac{3}{2})\]\[\frac{8}{3} - \frac{3}{2}\] Now find a common denominator (a number for the bottom that can be divided by both the current denominators): \[\frac{16}{6} - \frac{9}{6} = \frac{7}{6} = 1\frac{1}{6}\]

OpenStudy (anonymous):

thank you soooo much for your help!

OpenStudy (anonymous):

You're welcome!

hero (hero):

There's a more direct method that doesn't involve continually converting back and forth between mixed and improper.

OpenStudy (anonymous):

how do you do the direct method

hero (hero):

\[ \space\space\space\space\space2\frac{2}{3} \\-1\frac{1}{2} \\\text{_______} \] is equivalent to: \[ \space\space\space\space\space2\frac{4}{6} \\-1\frac{3}{6} \\\text{_______} \\\space\space\space\space\space{1\frac{1}{6}} \]

hero (hero):

^You thought this wasn't possible, but it is.

hero (hero):

Real funny @nincompoop

hero (hero):

No, I was referring to the form of it, not "changing to improper fractions".

hero (hero):

She hasn't seen it yet because she never came back online.

OpenStudy (anonymous):

@Hero is right, the direct method is easier in this particular case because there's no carrying involved. If the problem had been 2 1/2 - 1 2/3, however, it would have been more difficult. A good compromise is to convert the fractional part to a common denominator first, and if the one you're subtracting has a smaller numerator than the one your subtracting from they just do the math. If not, take my approach.